# Twisted character formula for toral supercuspidal representations

### v1

@@ 6.4.2. The case where $\alpha_{\mathrm{res}}$ is symmetric
as described in Lemma A.5 (each $\Gamma_{\mathbb{F}_{p}}$-orbit is symmetric). Note that $\epsilon([\eta]^{m_{\alpha}})=1$ for some $\epsilon\in P(V_{\Sigma\alpha},T)$ if and only if $\beta=\pm 1$, which is equivalent to that $\delta=1$. (This is furthermore equivalent to that $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}\neq 0$.) In this case, there exists a unique $\Gamma_{\mathbb{F}_{p}}$-orbit of such weights, hence the number $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})$ is given by r+s1r+s-1. Otherwise, we have $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})=r+s$. Since $\epsilon(-)^{\frac{1+q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{1}$ (and the same is true for $(k^{\prime}_{j})^{1}$), we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}\cdot\prod_{j=1}^{s}\operatorname{sgn}_{(k^{\prime}_{j})^{1}}$. In summary, we get
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{r+s-n_{\alpha}(\eta)}\cdot\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}(x_{i})\cdot\prod_{j=1}^{s}\operatorname{sgn}_{{k^{\prime}_{j}}^{1}}(x^{\prime}_{j})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta):=1$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta):=0$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}=0$.
@@
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=-1$, then we can apply the discussion in the case (1) to find a subtorus of $\operatorname{Sp}(V_{\Sigma\alpha})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=1$ and $f_{\alpha}^{\circ}$ is odd, then we can apply the discussion in the case (2) to find a subtorus of $\operatorname{Sp}(V_{\Sigma\alpha})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=1$ and $f_{\alpha}^{\circ}$ is even, then we repeat the discussion so far by furthermore replacing $\beta^{\frac{1}{2}}$ and $f_{\alpha}^{\circ}$ with $\beta^{\frac{1}{4}}$ and $\frac{f_{\alpha}^{\circ}}{2}$. Repeating this procedure, we can finally find a subtorus of $\operatorname{Sp}(V_{\Sigma\alpha})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
By applying the same consideration to $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{-\beta\}^{\frac{1}{f_{\alpha}}}$, in total, we can find an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ with an element whose multi-set of eigenvalues is the same as $\eta]^{m_{\alpha}}$.
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{k_{\alpha}^{\times}}(\beta)\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\beta)\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(-\beta)\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta)=1$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta)=0$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}=0$. By noting that
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
We furthermore note that, in the former case ($\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta)=-1$), the multi-set of eigenvalues $\mathcal{E}_{\mathbb{F}_{p}}([\eta]^{m_{\alpha}}\mid k_{\alpha}\oplus k_{\alpha})$ cannot contain $1$. In other words, $|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$ is in fact trivial. Keeping this in mind, we can write the result as
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$
@@
This means that the semisimple conjugacy class of the element $\eta]^{m_{\alpha}}\in\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$ does not depend on the choice of $\eta$. In other words, $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ is a constant determined only by $\alpha$.
@@
It is possible to explicate $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ by the same argument as in the unramified case, but we do not insist on it in this paper.
@@ 6.5. Twisted characters of Weil representations: symmetric roots
We next compute $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ in the case where $\alpha\in\Xi_{\mathrm{sym}}$. Note that $\alpha$ must be unramified by Lemma 6.2. Recall that $V_{\Sigma\alpha}=V_{\Gamma\alpha}\cong V_{\alpha}$ and that $V_{\alpha}$ is a $1$-dimensional $k_{\alpha}$-vector spaces, which is identified with $k_{\alpha}$ by fixing nonzero elements $X_{\alpha}\in V_{\alpha}$ (see Section 6.1).
@@
Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{\sigma_{\alpha}(\alpha)}$, the induced action of $\eta]^{m_{\alpha}}$ on $V_{\alpha}$ is $\varsigma_{\alpha}$-linear (note that $\eta]^{m_{\alpha}}$ is $k_{\alpha}$-linear on $V_{\Gamma\alpha}$):
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{\alpha}}$${V_{\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\Sigma\alpha})\cong\operatorname{Sp}(k_{\alpha})$. Note that, since $\eta]^{m_{\alpha}}$ preserves the symplectic form as described in Section 6.1, we must have
@@
To compute $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$, we follow the same strategy as in the case where $\alpha$ is asymmetric and $\alpha_{\mathrm{res}}$ is symmetric, i.e., appeal to Gérardin’s formula (Proposition A.3) by computing the eigenvalues of the action $\eta]^{m_{\alpha}}$ on $k_{\alpha}$. By Lemma A.10,
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ to be $\prod_{i=1}^{r}k_{i}^{1}$, then its element $(x_{i})_{i}$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\Sigma\alpha},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by
@@
as described in Lemma A.5 (note that each $\Gamma_{\mathbb{F}_{p}}$-orbit is symmetric). Note that $\epsilon([\eta]^{m_{\alpha}})=1$ for some $\epsilon\in P(V_{\Sigma\alpha},T)$ if and only if $\beta=1$. (This is furthermore equivalent to that $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}\neq 0$.) In this case, there exists a unique $\Gamma_{\mathbb{F}_{p}}$-orbit of such weights, hence the number $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})$ is given by r1r-1. Otherwise, we have $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})=r$. Since $\epsilon(-)^{\frac{1+q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{1}$, we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}$. In summary, we get
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{r-n_{\alpha}(\eta)}\cdot\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}(x_{i})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta)=1$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta)=0$ if $V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}=0$.
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{1-n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\eta_{\alpha})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
This means that, similarly to the case where $\alpha$ is asymmetric and $\alpha_{\mathrm{res}}$ is symmetric ramified, the semisimple conjugacy class of the element $\eta]^{m_{\alpha}}\in\operatorname{Sp}(k_{\alpha})$ does not depend on $\eta$. In other words, $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ is a constant determined only by $\alpha$.

### v2

@@ 6.4.2. The case where $\alpha_{\mathrm{res}}$ is symmetric
as described in Lemma A.5 (each $\Gamma_{\mathbb{F}_{p}}$-orbit is symmetric). Note that $\epsilon([\eta]^{m_{\alpha}})=1$ for some $\epsilon\in P(V_{\ddot{\alpha}},T)$ if and only if $\beta=\pm 1$, which is equivalent to that $\delta=1$. (This is furthermore equivalent to that $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}\neq 0$.) In this case, there exists a unique $\Gamma_{\mathbb{F}_{p}}$-orbit of such weights, hence the number $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})$ is given by r+s1r+s-1. Otherwise, we have $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})=r+s$. Since $\epsilon(-)^{\frac{1+q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{1}$ (and the same is true for $(k^{\prime}_{j})^{1}$), we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}\cdot\prod_{j=1}^{s}\operatorname{sgn}_{(k^{\prime}_{j})^{1}}$. In summary, we get
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{r+s-n_{\alpha}(\eta)}\cdot\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}(x_{i})\cdot\prod_{j=1}^{s}\operatorname{sgn}_{{k^{\prime}_{j}}^{1}}(x^{\prime}_{j})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta):=1$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta):=0$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}=0$.
@@
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=-1$, then we can apply the discussion in the case (1) to find a subtorus of $\operatorname{Sp}(V_{\ddot{\alpha}})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=1$ and $f_{\alpha}^{\circ}$ is odd, then we can apply the discussion in the case (2) to find a subtorus of $\operatorname{Sp}(V_{\ddot{\alpha}})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
    If $\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta^{\frac{1}{2}})=1$ and $f_{\alpha}^{\circ}$ is even, then we repeat the discussion so far by furthermore replacing $\beta^{\frac{1}{2}}$ and $f_{\alpha}^{\circ}$ with $\beta^{\frac{1}{4}}$ and $\frac{f_{\alpha}^{\circ}}{2}$. Repeating this procedure, we can finally find a subtorus of $\operatorname{Sp}(V_{\ddot{\alpha}})$ with an element whose multi-set of eigenvalues is $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{\beta\}^{\frac{1}{f_{\alpha}}}$.
By applying the same consideration to $\operatorname{Gal}(k_{\alpha_{\mathrm{res}}}/\mathbb{F}_{p})\cdot\{-\beta\}^{\frac{1}{f_{\alpha}}}$, in total, we can find an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ with an element whose multi-set of eigenvalues is the same as $\eta]^{m_{\alpha}}$.
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{k_{\alpha}^{\times}}(\beta)\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\beta)\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(-\beta)\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta)=1$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta)=0$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}=0$. By noting that
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
We furthermore note that, in the former case ($\operatorname{Nr}_{k_{\alpha_{\mathrm{res}}}/k_{\pm\alpha_{\mathrm{res}}}}(\beta)=-1$), the multi-set of eigenvalues $\mathcal{E}_{\mathbb{F}_{p}}([\eta]^{m_{\alpha}}\mid k_{\alpha}\oplus k_{\alpha})$ cannot contain $1$. In other words, $|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$ is in fact trivial. Keeping this in mind, we can write the result as
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{-\alpha}C)\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$
@@
This means that the semisimple conjugacy class of the element $\eta]^{m_{\alpha}}\in\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$ does not depend on the choice of $\eta$. In other words, $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ is a constant determined only by $\alpha$.
@@
It is possible to explicate $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ by the same argument as in the unramified case, but we do not insist on it in this paper.
@@ 6.5. Twisted characters of Weil representations: symmetric roots
We next compute $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ in the case where $\alpha\in\Xi_{\mathrm{sym}}$. Note that $\alpha$ must be unramified by Lemma 6.2. Recall that $V_{\ddot{\alpha}}=V_{\dot{\alpha}}\cong V_{\alpha}$ and that $V_{\alpha}$ is a $1$-dimensional $k_{\alpha}$-vector spaces, which is identified with $k_{\alpha}$ by fixing nonzero elements $X_{\alpha}\in V_{\alpha}$ (see Section 6.1).
@@
Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{\sigma_{\alpha}(\alpha)}$, the induced action of $\eta]^{m_{\alpha}}$ on $V_{\alpha}$ is $\varsigma_{\alpha}$-linear (note that $\eta]^{m_{\alpha}}$ is $k_{\alpha}$-linear on $V_{\dot{\alpha}}$):
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{\alpha}}$${V_{\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\ddot{\alpha}})\cong\operatorname{Sp}(k_{\alpha})$. Note that, since $\eta]^{m_{\alpha}}$ preserves the symplectic form as described in Section 6.1, we must have
@@
To compute $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$, we follow the same strategy as in the case where $\alpha$ is asymmetric and $\alpha_{\mathrm{res}}$ is symmetric, i.e., appeal to Gérardin’s formula (Proposition A.3) by computing the eigenvalues of the action $\eta]^{m_{\alpha}}$ on $k_{\alpha}$. By Lemma A.10,
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ to be $\prod_{i=1}^{r}k_{i}^{1}$, then its element $(x_{i})_{i}$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\ddot{\alpha}},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by
@@
as described in Lemma A.5 (note that each $\Gamma_{\mathbb{F}_{p}}$-orbit is symmetric). Note that $\epsilon([\eta]^{m_{\alpha}})=1$ for some $\epsilon\in P(V_{\ddot{\alpha}},T)$ if and only if $\beta=1$. (This is furthermore equivalent to that $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}\neq 0$.) In this case, there exists a unique $\Gamma_{\mathbb{F}_{p}}$-orbit of such weights, hence the number $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})$ is given by r1r-1. Otherwise, we have $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})=r$. Since $\epsilon(-)^{\frac{1+q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{1}$, we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}$. In summary, we get
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{r-n_{\alpha}(\eta)}\cdot\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{1}}(x_{i})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}},$
where $n_{\alpha}(\eta)=1$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}\neq 0$ and $n_{\alpha}(\eta)=0$ if $V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}=0$.
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{1-n_{\alpha}(\eta)}\cdot\operatorname{sgn}_{k_{\alpha}^{1}}(\eta_{\alpha})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
This means that, similarly to the case where $\alpha$ is asymmetric and $\alpha_{\mathrm{res}}$ is symmetric ramified, the semisimple conjugacy class of the element $\eta]^{m_{\alpha}}\in\operatorname{Sp}(k_{\alpha})$ does not depend on $\eta$. In other words, $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ is a constant determined only by $\alpha$.